# Peeling Context from Cause for Multimodal Molecular Property Prediction

### v1

@@ Abstract
Deep models are used for molecular property prediction, yet they are often hard to interpret and may rely on spurious context rather than causal structure, which degrades reliability under distribution shift and harms predictive performance. We introduce CLaP, Causal Layerwise Peeling, a framework which separates causal signal from context in a layerwise manner and integrates diverse graph representations of molecules. At each layer, a causal block performs a soft split into causal and trivial branches, fuses causal evidence across modalities, and progressively peels batch-coupled context to concentrate on label-relevant structure, thereby limiting shortcut signals and stabilizing layerwise refinement. Across four molecular benchmarks, CLaP reliably improves MAE, MSE, and $R^{2}$ over competitive baselines. We also obtain atom-level causal saliency maps that highlight substructures responsible for a prediction, providing actionable guidance for targeted molecular edits. Case studies confirm the accuracy of these maps and their alignment with chemical intuition. By peeling context from cause at every layer, the model delivers predictors that are accurate and interpretable for molecular design.
1 Introduction
Designing molecules with desired properties is a central goal in drug discovery and materials design (SanchezLengeling2018Inverse). Graph-based deep learning is effective for property prediction (Wu2018MoleculeNet; Hinton06; Bengio+chapter2007; goodfellow2016deep). However, models often exploit spurious correlations tied to datasets or batches (Geirhos2020Shortcut), which hurts reliability under distribution shift. They also provide limited substructure-level guidance (JimenezLuna2020XAI), reducing their value for design. These gaps motivate predictors that separate causal signal from contextual shortcuts and yield chemically meaningful attributions.
Prior work on invariant rationales (e.g., DIR, CAL) promotes causal subgraphs by enforcing invariance across constructed environments (DIR; CAL; CIGA). While influential, these approaches are largely designed for classification, rely on synthetic environments whose fidelity to real data is uncertain, and often treat causality as a one-shot selection problem. We take a different route and introduce CLaP, Causal Layerwise Peeling, a layerwise framework that peels context from cause. At each layer, a learnable splitter routes features into a causal branch and a trivial branch, progressively peeling batch-coupled context and concentrating on label-relevant structure. Rather than engineering synthetic environments (e.g., graph perturbations) or selecting a sparse subgraph with a predictor trained to be stable across them, we exploit the natural fluctuations of ordinary mini-batches during training as contextual variation that the model learns to ignore.
Our design is grounded in a batch-wise invariance principle. As batch composition changes during training, only sample intrinsic signal should keep a stable alignment with the label. We implement this with a depth dependent correlation target and a monotonicity regularizer. The causal readout increases its within batch Pearson correlation with the label from shallow to deep layers. A stop gradient residual objective trains the trivial branch on the remaining error. This enforces a clean division of labor and prevents shortcut leakage into the causal path. Building on this premise, we formulate a regression-ready objective that aligns the scalar causal readout with continuous labels and enforces non-decreasing correlation across depth. These design choices enable CLaP to peel context from cause without synthetic environments and to provide robust, interpretable predictions.
From a molecular-property standpoint, different representations usually capture complementary facets of the same molecule. We therefore fuse same-depth causal evidence from three views—2D topology (SMILES graphs), peptide notation (HELM (Zhang2012HELM)), and 3D geometry. A learned gating module assigns sample-specific weights, up-weighting the modality that carries batch-invariant signal at each layer. Gating weights yield atom-level causal saliency maps, and layerwise fusion peels away context to reveal the label-driving substructures (see Fig. 1).
@@ 1 Introduction
Figure 1: The original molecule in this illustrative example is iteratively “peeled” across layers. At shallow layers (layer 1), attribution underpeels and overweighs the carbon backbone. At the optimal layer, causal weights concentrate on the nitro group, yielding chemically plausible solubility drivers. Deeper peeling begins to smear attribution across the backbone, reflecting over-peeling in practice.
Experiments on small-molecule benchmarks (ESOL, FreeSolv, Lipo) and cyclic peptides (CycPeptMPDB) show CLaP achieves consistent gains in MAE, MSE, and $R^{2}$ over architecture-centric baselines (MAT, FP-GNN) and causality-oriented baselines (DIR, CAL) (MAT; Cai2022FPGNN; DIR; CAL). The atom-level saliency maps align with chemical intuition and, in case studies, pinpoint causal patterns and suggest targeted molecular edits.
@@
  • We propose CLaP, a multimodal, layerwise framework that splits features into causal and trivial branches at every layer and fuses same layer causal evidence across molecular views, reducing reliance on datasets or batches specific shortcuts, enhancing causality.
  • CLaP produces causal saliency maps that highlight label-critical substructures. In case studies, these maps match chemical intuition and provide guidance for molecular design.
  • Across diverse molecular property datasets, CLaP surpasses strong baselines. Its causal–trivial split reduces reliance on spurious context, improving robustness and accuracy.
2 Methodology
We first formalize a batch-wise invariance principle that provides the theoretical basis for our supervised causal objective. We then instantiate it with a layerwise causal–trivial architecture that fuses modalities and splits features per layer to enable both causal alignment and interpretability.
2.1 Batch-wise Invariance
During training, each molecule is repeatedly sampled into different mini-batches across epochs, and thus encounters varying batch contexts defined by the co-occurring samples and their label statistics. A reliable predictor should maintain consistent alignment with the ground-truth label regardless of such batch fluctuations. This observation motivates a batch-wise invariance principle: only sample-intrinsic signal can sustain stable correlation with the label under changing batch contexts.
Figure 2: Overview of the causal layerwise peeling architecture for $L=2$ layers.
@@ 2.1 Batch-wise Invariance
Index mini-batches by $e\in\mathcal{E}$ over random shuffles. Assume an additive view of the label
$y\;=\;\theta\,c(x)\;+\;\eta,\qquad\mathbb{E}[\eta\mid x]=0,$ (1)
where $c(x)$ is a sample-intrinsic, label-relevant component. Let $q(x,e)$ denote context features whose association with $y$ may depend on the batch $e$. Consider a local linearization of a layer’s scalar predictor with scalars $a,b$.
s(x,e)=ac(x)+bq(x,e),s(x,e)\;=\;a\,c(x)\;+\;b\,q(x,e), (2)
Batchwise correlation.
For any batch $e$ of size $B_{e}$, define batch-centered variables
$\widetilde{u}_{e}\;=\;u\;-\;\tfrac{1}{B_{e}}(\mathbf{1}^{\top}u)\,\mathbf{1},$
and the within-batch Pearson correlation
$\mathrm{Corr}_{e}(u,v)\;=\;\frac{\langle\widetilde{u}_{e},\,\widetilde{v}_{e}\rangle}{\|\widetilde{u}_{e}\|_{2}\,\|\widetilde{v}_{e}\|_{2}+\varepsilon},\qquad\varepsilon>0\ \text{small}.$ (3)
@@
Batch-wise invariance requires
$\mathrm{Corr}_{e}\!\big(s(x,e),\,y\big)\;\approx\;\rho\quad\text{for all }e\in\mathcal{E},\ \text{fixed}\ \rho\in[-1,1]\ .$ (4)
i.e., small dispersion of $\mathrm{Corr}_{e}(s,y)$ across changing batches. A convenient surrogate is
$\min_{a,b}\;\sum_{e\in\mathcal{E}}\Big(\mathrm{Corr}_{e}\!\big(s(x,e),y\big)-\rho\Big)^{2}.$ (5)
@@
    $\mathrm{Corr}_{e}\!\big(c(x),y\big)=\kappa$ is constant in $e$ (sample-intrinsic stability).
@@
Then any predictor of the form equation 2 that satisfies the batch-wise invariance requirement equation 4 for all $e\in\mathcal{E}$ must have $b=0$. Equivalently, the batch-invariant solution discards $q$ and retains $c$, and $\mathrm{Corr}_{e}(s,y)=\mathrm{Corr}_{e}(c,y)=\kappa$ for all $e$. A proof for invariance is provided in Appendix A.1.
2.2 Layer-wise Causal–Trivial Peeling
Our framework integrates three molecular views—2D graphs (SM), HELM notations (PE), and 3D geometries (GE). At each layer, causal blocks softly split features into causal and trivial paths. Causal signals are fused across modalities and aligned with the label via a depth-wise correlation schedule, while residual context is peeled and accumulated with stop-gradient. The final prediction combines the last causal readout and the accumulated residual (see Fig. 2).
@@ 2.2 Layer-wise Causal–Trivial Peeling
From the modality-specific node embeddings produced by the backbone encoders, each block at depth $\ell\!\in\!\{1,\dots,L\}$ applies a learnable node-wise splitter that assigns every atom/group a causal gate $\alpha_{c}^{(\ell)}\!\in\![0,1]$. Its complement $\alpha_{t}^{(\ell)}\!=\!1-\alpha_{c}^{(\ell)}$ acts as the trivial gate. These gates induce two branches per modality: a causal representation ${\bm{z}}_{c}^{(\ell)}$ and a trivial representation ${\bm{z}}_{t}^{(\ell)}$. For each modality at depth $\ell$, we obtain causal vectors ${\bm{z}}_{\textit{SM}}^{(\ell)},\ {\bm{z}}_{\textit{PE}}^{(\ell)},\ {\bm{z}}_{\textit{GE}}^{(\ell)}\!\in\!\mathbb{R}^{D}$ from the current block, along with the trivial vectors ${\bm{z}}_{\textit{SM}}^{\prime(\ell)},\ {\bm{z}}_{\textit{PE}}^{\prime(\ell)},\ {\bm{z}}_{\textit{GE}}^{\prime(\ell)}\!\in\!\mathbb{R}^{D}$. A gating network $g_{F}(\cdot)$ produces sample-specific soft weights $w^{(\ell)}$ over modalities via a softmax, and these weights are used to form the fused causal representation ${\bm{z}}_{\textit{fuse}}^{(\ell)}$ (and the fused trivial representation $\mathbf{z}_{\mathrm{\textit{fuse}}}^{\prime(\ell)}$) at depth $\ell$.
@@
Scalar readouts use projection heads. Applying causal head ${\bm{u}}^{(\ell)}$ to $\mathbf{z}_{\textit{fuse}}^{(\ell)}$ yields ${\bm{c}}^{(\ell)}$, while applying trivial head ${\bm{v}}^{(\ell)}$ to $\mathbf{z}_{\textit{fuse}}^{\prime(\ell)}$ yields ${\bm{t}}^{(\ell)}$. The causal branch is forwarded to depth $\ell{+}1$ (yielding ${\bm{c}}^{(\ell+1)}$) while an additional trivial component $t^{(\ell+1)}$ is peeled. This recursion continues until depth $L$.
After block $L$, the per-batch scalars are stacked across layers to enable supervision and analysis.
@@
The causal branch is designed to focus, layerwise, on the label-predictive signal. At each depth $\ell$, we align the fused causal scalar ${\bm{c}}^{(\ell)}$ with the batch labels ${\bm{y}}\!\in\!\mathbb{R}^{B}$ by constraining their Pearson correlation to follow a depth-dependent target schedule.
We first center the variables and then calculate the mini-batch Pearson correlation at depth $\ell$ with a small stabilizer ${\epsilon}>0$.
@@
$\begin{gathered}\begin{aligned} \widetilde{{\bm{c}}}^{(\ell)}&={\bm{c}}^{(\ell)}-\tfrac{1}{B}({\bm{1}}^{\top}{\bm{c}}^{(\ell)})\,{\bm{1}}\qquad\widetilde{{\bm{y}}}&={\bm{y}}-\tfrac{1}{B}({\bm{1}}^{\top}{\bm{y}})\,{\bm{1}},\end{aligned}\\[2.0pt] \mathrm{corr}_{\ell}=\frac{\langle\widetilde{{\bm{c}}}^{(\ell)},\,\widetilde{{\bm{y}}}\rangle}{\|\widetilde{{\bm{c}}}^{(\ell)}\|_{2}\,\|\widetilde{{\bm{y}}}\|_{2}+{\epsilon}}\in[-1,1].\end{gathered}$ (6)
@@
We set a linearly increasing target correlation schedule so that deeper layers are encouraged to capture progressively finer, label-critical variation:
$\rho_{\ell}=\rho_{\min}+\frac{\ell-1}{L-1}\,(\rho_{\max}-\rho_{\min}),\qquad\ell=1,\dots,L.$ (7)
Then penalize deviations between the batchwise Pearson correlation and its target,
$\mathcal{L}_{\mathrm{corr}}=\frac{1}{L}\sum_{\ell=1}^{L}\big(\mathrm{corr}_{\ell}-\rho_{\ell}\big)^{2},$ (8)
@@
The correlation loss guides each layer toward the target depth-wise trend and concentrates causal signal while filtering out irrelevant batch-specific context noise.
@@
To stabilize the hierarchy and prevent causal regression from $\ell$ to $\ell{+}1$, we further apply a monotonicity regularizer, encouraging correlations to be non-decreasing from layer to layer via a hinge penalty with margin $\gamma\!\geq\!0$:
$\mathcal{L}_{\mathrm{mono}}=\frac{1}{L-1}\sum_{\ell=1}^{L-1}\max\!\Big\{0,\ \mathrm{corr}_{\ell}-\mathrm{corr}_{\ell+1}+\gamma\Big\}.$ (9)
@@
Together, equation 8–equation 9 encourage a depth-wise progression. Shallow layers capture coarse, high-variance patterns, and deeper layers increasingly model causal patterns.
Trivial branch: residual learning.
@@
The trivial branch isolates context and irrelevant variation that are not central to prediction. By removing these factors from the causal branch, the causal signal becomes clearer and more discriminative. At depth $\ell$, each block yields a trivial scalar ${\bm{t}}^{(\ell)}$. We aggregate these across layers, while retaining only the final causal scalar as the causal output:
$t=\sum_{\ell=1}^{L}{\bm{t}}^{(\ell)}\in\mathbb{R}^{B},\qquad y_{\mathrm{c}}^{\star}={\bm{c}}^{(L)}.$ (10)
@@
To maintain a clear division of labor, the trivial branch is trained on the residual between the ground-truth label $\bm{y}$} and the final causal output, with gradients stopped through $y_{\mathrm{c}}^{\star}$ to avoid contaminating the causal pathway:
$\mathcal{L}_{\mathrm{triv}}=\mathrm{MSE}\!\Big(t,\ {\bm{y}}-\mathrm{stopgrad}(y_{\mathrm{c}}^{\star})\Big).$ (11)
@@
This design assigns the stable, generalizable component to the causal path, whereas the trivial path captures the remaining residual variation, calibrating $y_{\mathrm{c}}^{\star}$ by absorbing context-specific noise that should not be attributed to the causal branch.
@@
The model predicts by additively combining the depth-$L$ causal readout with the accumulated trivial correction:
$\widehat{y}=y_{\mathrm{c}}^{\star}+t,\qquad\mathcal{L}_{\mathrm{pred}}=\operatorname{MSE}(\widehat{y},\,{\bm{y}}).$ (12)
@@
We regularize the causal branch toward batch-invariant, depth-increasing alignment with the label, while training the trivial branch on the residual (with stop–gradient on $y_{\mathrm{c}}^{\star}$). The full objective is
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{pred}}+\underbrace{\lambda_{\mathrm{caus}}\,\mathcal{L}_{\mathrm{corr}}}_{\text{causal alignment}}+\underbrace{\lambda_{\mathrm{mono}}\,\mathcal{L}_{\mathrm{mono}}}_{\text{depth monotonicity}}+\underbrace{\lambda_{\mathrm{unif}}\,\mathcal{L}_{\mathrm{triv}}}_{\text{residual fit}},\quad\lambda_{\mathrm{caus}},\lambda_{\mathrm{mono}},\lambda_{\mathrm{unif}}\geq 0.$
3 Experiments
Figure 3: Performance comparison across four molecular property datasets.
3.1 Experimental Settings
Datasets. We evaluate CLaP on four publicly available benchmark datasets: ESOL, a widely used dataset of aqueous solubility values; FreeSolv, reporting experimentally measured hydration free energies of small molecules; and Lipo, containing curated standard octanol/water distribution coefficients; CycPeptMPDB, which measures passive cell permeability of cyclic peptides.

### v2

# Peeling Context from Cause for Molecular Property Prediction

@@ Abstract.
Deep models are used for molecular property prediction, yet they are often hard to interpret and may rely on spurious context rather than causal structure, degrading reliability under distribution shift and harming predictive performance. We introduce CLaP, Causal Layerwise Peeling, a framework that separates causal signal from context in a layerwise manner and integrates diverse graph representations of molecules. At each layer, a causal block performs a soft split into causal and context branches, fuses causal evidence across modalities, and peels batch-coupled context to concentrate on label-relevant structure, limiting shortcut signals and stabilizing layerwise refinement. Across nine molecular benchmarks, including OOD settings, CLaP reliably reduces MAE and MSE relative to competitive baselines. Beyond predictive performance, we conduct targeted causal evaluations under controlled environment shifts. These analyses provide empirical evidence of clear separation between environment-dependent context and causal signals. As further evidence, we also analyze atom-level causal saliency maps and counterfactual outcomes, which highlight property-relevant substructures and show consistent alignment with chemical intuition.
molecular property prediction; multimodal graph learning; causal inference
$c$s: Applied computing Chemistryccs: Computing methodologies Machine learning
Refer to caption Illustration of layer-wise causal peeling on a molecule, showing attribution shifts from peripheral groups to the backbone across layers.
@@
Figure 1. The original molecule in this illustrative example is iteratively “peeled” across layers. At shallow layers (layer 1), attribution underpeels the carbon backbone. At the optimal layer, causal weights concentrate on the nitro group, yielding chemically plausible solubility drivers. Deeper peeling begins to smear attribution across the backbone, reflecting over-peeling.
1. Introduction
Designing molecules with desired properties is a central goal in drug discovery and materials design (Sanchez-Lengeling and Aspuru-Guzik, 2018). Graph-based deep learning is effective for property prediction (Wu et al., 2018; Goodfellow et al., 2016). However, models often inherently exploit spurious correlations tied to datasets or batches (Geirhos et al., 2020), which hurts reliability under distribution shift (Arjovsky et al., 2019; Li et al., 2025a, 2024; An et al., 2025). They also provide limited interpretable substructure-level guidance (Jiménez-Luna et al., 2020), reducing their value for design. These gaps motivate the development of predictors that separate causal signal from contextual shortcuts and yield chemically meaningful attributions (Sui et al., 2022; Yin et al., 2025).
Figure 2. Overview of the causal layerwise peeling architecture for $L=2$ layers.
Prior work on invariant rationales promotes causal subgraphs by enforcing invariance across constructed environments (Wu et al., 2024; Wang et al., 2024b; Wu et al., 2022; Sui et al., 2022; Chen et al., 2022; Peters et al., 2016). While influential, these approaches are largely designed for classification, rely on synthetic environments whose fidelity to real data is uncertain, and often treat causality as a one-shot selection problem. We take a different route and introduce CLaP, Causal Layerwise Peeling, a layerwise framework that peels context from cause. At each layer, a learnable splitter routes features into a causal branch and a context branch, progressively peeling batch-coupled context and concentrating on label-relevant structure. Rather than engineering synthetic environments such as varying graph perturbations, or selecting a sparse subgraph with a predictor trained to be stable across them, we exploit the natural fluctuations of mini-batches as a source of contextual variation, which the model explicitly separates and absorbs through a dedicated context branch.
Our design is grounded in a batch-wise invariance principle. As batch composition changes during training, only sample intrinsic signal should keep a stable alignment with the label. We implement this with a depth-dependent correlation target and a monotonicity regularizer. The causal readout increases its within-batch Pearson correlation with the label from shallow to deep layers. A residual objective trains the context branch on the remaining error. This enforces a clean division of labor and prevents shortcut leakage into the causal path. Building on this premise, we formulate a regression-ready objective that aligns the scalar causal readout with continuous labels and enforces non-decreasing correlation across depth. These designs enable CLaP to peel context from cause without synthetic environments and to provide robust, interpretable predictions.
From a molecular-property standpoint, different representations usually capture complementary facets of the same molecule (Xu et al., 2024; Zhang et al., 2024; Luo and Deng, 2025). We therefore fuse causal evidence from three views—2D topology, peptide notation (Zhang et al., 2012), and 3D geometry (Schütt et al., 2017). A gating module assigns sample-specific weights, up-weighting the modality that carries batch-invariant signal at each layer. Gating weights yield atom-level causal saliency maps, and layerwise fusion peels away context to reveal label-driving substructures (see Fig. 1).
Experiments on small-molecule benchmarks (ESOL, FreeSolv, Lipo) (Delaney, 2004; Mobley and Guthrie, 2014; Hersey, 2015) and cyclic peptides (CycPeptMPDB) (Li et al., 2023), as well as OOD evaluations on Half_Life_Obach, Solubility, LD50_Zhu, Hepatocyte, and Microsome (Huang et al., 2021), show that CLaP achieves the most consistent improvements in MAE and MSE over both architecture-centric baselines (MAT, FP-GNN, MolFCL, GSL) (Maziarka et al., 2019; Cai et al., 2022; Tang et al., 2025; Zhao et al., 2024) and causality-oriented baselines (DIR, CAL, CGR) (Wu et al., 2022; Sui et al., 2022; Yin et al., 2025). Beyond predictive improvements, we explicitly evaluate whether the learned representations exhibit the intended causal behavior. We design controlled environment shifts at test time, including structural aggregation and re-batching interventions, to examine how the causal and context branches respond. These evaluations indicate that CLaP tends to isolate signals that remain stable across different environments. In addition, atom-level causal saliency maps naturally induced by the model’s causal gating mechanism are consistent with known structure–property relationships and, through case studies and counterfactual analyses, help identify substructures that are likely to causally influence the target properties.
@@ 1. Introduction
  • We propose CLaP, a multimodal, layerwise framework that separates causal and context representations and fuses causal evidence across molecular views, achieving empirically consistent performance improvements over strong baselines while improving prediction accuracy and robustness.
  • We evaluate the causal behavior of learned representations under controlled environment shifts, including scaffold-level aggregation and re-batching, to examine signal stability and the separation between causal and context branches.
  • CLaP yields atom-level causal saliency maps induced by its gating mechanism that highlight property-relevant substructures consistent with structure–property relationships.
2. Methodology
We first formalize a batch-wise invariance principle that provides the theoretical basis for our supervised causal objective. We then instantiate it with a layerwise causal–context architecture.
2.1. Batch-wise Invariance
During training, each molecule is repeatedly sampled into different mini-batches across epochs, and thus encounters varying batch contexts defined by co-occurring samples and label statistics. A reliable predictor should maintain consistent alignment with the ground-truth label regardless of such batch fluctuations. This observation motivates a batch-wise invariance principle: only a sample-intrinsic signal sustains stable correlation with labels under batch contexts.
@@ 2.1. Batch-wise Invariance
Index mini-batches by $e\in\mathcal{E}$ over random shuffles. Assume an additive view of the prediction $s(x,e)$ where $c(x)$ is a sample-intrinsic component. Let $q(x,e)$ denote context features whose association with $y$ may depend on the batch $e$, and consider a linearization of the layer’s scalar predictor with coefficients $a,b$.
(1) s(x,e)=ac(x)+bq(x,e),s(x,e)\;=\;a\,c(x)\;+\;b\,q(x,e),
@@
For any batch $e$ of size $B_{e}$, define batch-centered variables ${s}_{e}$ and the within-batch Pearson correlation. Batch-wise invariance requires
(2) $\mathrm{Corr}_{e}\!\big(s(x,e),\,y\big)\;\approx\;\rho\quad\text{for all }e\in\mathcal{E},\ \text{fixed}\ \rho\in[-1,1]\ .$
i.e., small dispersion of $\mathrm{Corr}_{e}(s,y)$ across changing batches.
(3) $\min_{a,b}\;\sum_{e\in\mathcal{E}}\Big(\mathrm{Corr}_{e}\!\big(s(x,e),y\big)-\rho\Big)^{2}.$
@@
    $\mathrm{Corr}_{e}\!\big(c(x),y\big)=\kappa$ is constant in $e$ (sample-intrinsic).
@@
Then any predictor of the form equation 1 that satisfies the batch-wise invariance requirement equation 2 for all $e\in\mathcal{E}$ must necessarily have $b=0$. Equivalently, the batch-invariant solution thus only discards $q$ and retains $c$, and $\mathrm{Corr}_{e}(s,y)=\mathrm{Corr}_{e}(c,y)=\kappa$ for all $e$. A proof for invariance is provided in Appendix A.1.
2.2. Layer-wise Causal–context Peeling
Our framework integrates three molecular views—2D graphs (SM), HELM notations (PE), and 3D geometries (GE). At each layer, causal blocks softly split features into causal and context paths. Causal signals are fused across modalities and aligned with the label via a depth-wise correlation schedule, while residual context is peeled and accumulated with stop-gradient. The final prediction combines the last causal readout and the accumulated residual (see Fig. 2).
@@ 2.2. Layer-wise Causal–context Peeling
From the modality-specific node embeddings produced by the backbone encoders, each block at depth $\ell\!\in\!\{1,\dots,L\}$ applies a learnable node-wise splitter that assigns every atom/group a causal gate $\alpha_{c}^{(\ell)}\!\in\![0,1]$. Its complement $\alpha_{t}^{(\ell)}\!=\!1-\alpha_{c}^{(\ell)}$ acts as the context gate. These gates induce two branches per modality: a causal representation ${\bm{z}}_{c}^{(\ell)}$ and a context representation ${\bm{z}}_{t}^{(\ell)}$. For each modality at depth $\ell$, we obtain causal vectors ${\bm{z}}_{\textit{SM}}^{(\ell)},\ {\bm{z}}_{\textit{PE}}^{(\ell)},\ {\bm{z}}_{\textit{GE}}^{(\ell)}\!\in\!\mathbb{R}^{D}$ from the current block, along with the context vectors ${\bm{z}}_{\textit{SM}}^{\prime(\ell)},\ {\bm{z}}_{\textit{PE}}^{\prime(\ell)},\\$ ${\bm{z}}_{\textit{GE}}^{\prime(\ell)}\!\in\!\mathbb{R}^{D}$. A gating network $g_{F}(\cdot)$ produces sample-specific soft weights $w^{(\ell)}$ over modalities via a softmax, and these weights are used to form the fused causal representation ${\bm{z}}_{\textit{fuse}}^{(\ell)}$ (and the fused context representation $\mathbf{z}_{\mathrm{\textit{fuse}}}^{\prime(\ell)}$) at depth $\ell$.
@@
Scalar readouts use projection heads. Applying causal head ${\bm{u}}^{(\ell)}$ to $\mathbf{z}_{\textit{fuse}}^{(\ell)}$ yields ${\bm{c}}^{(\ell)}$, while applying context head ${\bm{v}}^{(\ell)}$ to $\mathbf{z}_{\textit{fuse}}^{\prime(\ell)}$ yields ${\bm{t}}^{(\ell)}$. The causal branch is forwarded to depth $\ell{+}1$ (yielding ${\bm{c}}^{(\ell+1)}$) while an additional context component $t^{(\ell+1)}$ is peeled. This recursion continues until depth $L$. After block $L$, the per-batch scalars are stacked across layers to enable supervision and analysis.
@@
The causal branch is designed to focus, layerwise, on the label-predictive signal. At each depth $\ell$, we align the fused causal scalar ${\bm{c}}^{(\ell)}$ with the batch labels ${\bm{y}}\!\in\!\mathbb{R}^{B}$ by explicitly constraining their Pearson correlation to follow a depth-dependent target schedule in a supervised manner.
@@
(4) $\begin{gathered}\mathrm{corr}_{\ell}=\frac{\langle{\bm{c}}^{(\ell)},\,{{\bm{y}}}\rangle}{\|{{\bm{c}}}^{(\ell)}\|_{2}\,\|{{\bm{y}}}\|_{2}+{\epsilon}}\in[-1,1].\end{gathered}$
@@
We set a linearly increasing target correlation schedule $\rho_{\ell}$ from $\rho_{\min}$ to $\rho_{\max}$ so that deeper layers are encouraged to capture progressively finer, label-critical variation. Then penalize deviations between the batchwise Pearson correlation and its target,
(5) $\mathcal{L}_{\mathrm{corr}}=\frac{1}{L}\sum_{\ell=1}^{L}\big(\mathrm{corr}_{\ell}-\rho_{\ell}\big)^{2},$
@@
The correlation loss guides each layer toward the target depth-wise trend and concentrates causal signal while filtering out irrelevant batch-specific context noise in a principled way.
@@
To stabilize the hierarchy and ensure consistent refinement of causal regression from $\ell$ to $\ell{+}1$, we further apply a monotonicity regularizer, encouraging correlations to be non-decreasing from layer to layer via a hinge penalty with margin $\gamma\!\geq\!0$:
(6) $\mathcal{L}_{\mathrm{mono}}=\frac{1}{L-1}\sum_{\ell=1}^{L-1}\max\!\Big\{0,\ \mathrm{corr}_{\ell}-\mathrm{corr}_{\ell+1}+\gamma\Big\}.$
@@
Together, equation 5–equation 6 encourage a depth-wise progression of representation. Shallow layers capture coarse, high-variance patterns, and deeper layers increasingly model causal patterns.
Table 1. Comparison across nine benchmarks and ablation results. Best per-dataset/metric in bold. Lower is better.
Method ESOL FreeSolv Lipo CycPeptMPDB Half Life Hepatocyte Microsome Solubility Toxicity
MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
[…]
GSL 1.0032 1.2566 2.3476 4.0343 0.9007 1.1478 0.6590 0.8127 0.9365 1.3304 1.1045 1.3201 1.0723 1.2428 1.5883 1.9713 0.7414 0.9073
[…]
CGR 1.5853 4.0291 2.7268 11.4813 0.9670 1.5269 0.5792 0.5578 1.0360 1.6673 1.2371 1.9796 1.2544 1.9550 1.6741 4.5156 0.7794 1.0112
[…]
MolFCL 0.4660 0.4740 1.0300 1.9890 0.4330 0.3670 0.3690 0.2440 0.9210 1.5000 1.1160 1.7660 0.9430 1.3580 0.6990 0.9060 0.7530 1.0920
[…]
w/o correlation 0.5078 0.4767 0.8779 1.3742 0.5562 0.5122 0.3215 0.1866 0.8222 1.1382 1.0230 1.5725 0.8051 1.0930 0.7835 1.1295 0.5250 0.5087
[…]
Context branch: context learning.
@@
The context branch isolates environment- and batch-dependent variation that may contribute to prediction but is not stable across environments in practice. By removing these factors from the causal branch, the causal signal becomes clearer and more discriminative. At depth $\ell$, each block yields a context scalar ${\bm{t}}^{(\ell)}$. We aggregate these across layers, while retaining only the final causal scalar as the causal output:
(7) $t=\sum_{\ell=1}^{L}{\bm{t}}^{(\ell)}\in\mathbb{R}^{B},\qquad y_{\mathrm{c}}^{\star}={\bm{c}}^{(L)}.$
@@
To maintain a clear division of labor, the context branch is trained on the residual between label $\bm{y}$} and final causal output, with gradients stopped through $y_{\mathrm{c}}^{\star}$ to avoid contaminating the causal pathway:
(8) $\mathcal{L}_{\mathrm{ctx}}=\mathrm{MSE}\!\Big(t,\ {\bm{y}}-\mathrm{stopgrad}(y_{\mathrm{c}}^{\star})\Big).$
@@
This design assigns the stable, generalizable component to the causal path, while the context path captures the remaining residual variation, calibrating $y_{\mathrm{c}}^{\star}$ by absorbing context-specific noise, such as molecular scaffolds or batch-level statistical effects.
@@
The model predicts by additively combining depth-$L$ causal readout with the accumulated context:
(9) $\widehat{y}=y_{\mathrm{c}}^{\star}+t,\qquad\mathcal{L}_{\mathrm{pred}}=\operatorname{MSE}(\widehat{y},\,{\bm{y}}).$
@@
We regularize the causal branch toward batch-invariant, depth-increasing alignment with the label, while training the context branch on the residual. The overall objective function integrates these structural priors into a joint loss formulated as:
[…]
3. Experiments
3.1. Experimental Setup
Datasets. We evaluate CLaP on nine molecular property prediction datasets, categorized into two standardized evaluation settings: in-distribution (ID) and out-of-distribution (OOD). An overview of dataset statistics is provided in Table 2(a).
In the in-distribution setting, we apply random splits to four representative small-molecule and peptide benchmarks commonly used in prior work: ESOL (aqueous solubility) (Delaney, 2004), FreeSolv (hydration free energy) (Mobley and Guthrie, 2014), Lipo (octanol/water partition coefficient) (Hersey, 2015), and CycPeptMPDB (passive permeability of cyclic peptides) (Li et al., 2023).
To evaluate robustness under distribution shift, we adopt preset scaffold-based OOD splits on five additional datasets from TDC (Huang et al., 2021): Half_Life_Obach, Solubility, Toxicity, Hepatocyte, and Microsome, which cover absorption, excretion, and toxicity properties.